OpenStudy (annjewell1998):

Let f(x) = 9x - 4.5. Find the value of f(2). a 6.5 b 87.5 c 13.5 d 9 Okay so I know that you're supposed to insert 2 for x and solve I just don't know how to solve it from there. Please help?

Mehek (mehek14):

put 2 for x \(9*2-4.5\\9*2=?\)

OpenStudy (calculusxy):

You need to replace \(x\) for 2. \(f(x) = 9(2) - 4.5\)

OpenStudy (calculusxy):

Sorry wait. \(f(2) = 9(2) - 4.5\)

OpenStudy (annjewell1998):

So 9*2=18-4.5?

Mehek (mehek14):

yup

OpenStudy (calculusxy):

now you just need to evaluate \(18 - 4.5\)

OpenStudy (annjewell1998):

13.5. So 13.5 is the answer?

Mehek (mehek14):

yes

OpenStudy (calculusxy):

correct!

OpenStudy (annjewell1998):

Thanks you guys!

Mehek (mehek14):

yw :)